OpenStudy (anonymous):

Triangle ABC is congruent to triangle DEF. In triangle ABC, side AB measures 13, side BC measures x+13, and side CA measures 16. In triangle DEF, side DE measures 13, side EF measures 3x−20, and side FD measures 16. What equation would help you to solve for the side length of BC and EF? Explain your reasoning using complete sentences.

OpenStudy (anonymous):

Since the triangles are congruent, the sides and angles are the same. So AB=DE, BC=EF, and AC=DF Since BC=EF just substitute the given value for each side and solve for x. So you would have this: x+13=3x-20 2x=33 x=33/2 or 16.5 So BC = EF= 16.5

OpenStudy (anonymous):

Actually, BC = EF = 29.5